# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v1

@@ 1.2 The Low-Degree Framework
We must formally define what it means for such a polynomial to find an independent set in a hypergraph. We follow the notation and definitions from [Wei22], in which a rounding procedure is used to produce an independent set of $\mathcal{H}_{2}(n,p)$ based on the output $f(A)$. For brevity, we just state the following definition in this section for ordinary hypergraphs (the multipartite analog can be inferred as it is nearly identical).
@@
[…]
@@ 1.3 Main Results
[…]
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, the following hold.
@@
    $k_{j}\leq\gamma_{j}\,r\,n(1-\varepsilon)\left(\frac{\log d}{d\,r^{r}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@
    $k_{j}\,\geq\,\gamma_{j}\,r\,n(1+\varepsilon)\left(\frac{\log d}{d\,r^{r}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@ 1.4 Proof Overview
Let us first discuss the achievability result (Theorem 1.5). As observed by Wein in [Wei22], sparse random graphs are locally tree-like. In particular, the $s$-neighborhood of a vertex in $\mathcal{H}_{2}(n,d/n)$ can be well approximated by the Poisson Galton-Watson tree, a random graph model for generating trees. In [PZ21], Pal and Zhu generalized this model to the hypergraph setting, describing the $r$-uniform Poisson Galton-Watson Hypertree $\mathbb{T}_{d}^{GW}$ (see §2.2 for a formal description). We observe that for $p=d/\binom{n-1}{r-1}$, the hypergraph $\mathcal{H}_{r}(n,p)$ is locally $\mathbb{T}_{d}^{GW}$-like, which implies that the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding algorithm evaluated at the root of $\mathbb{T}_{d}^{GW}$. Our proof now follows in two steps: first, we describe a local algorithm for finding large independent sets in $\mathbb{T}_{d}^{GW}$ (which can be adapted to one on $\mathcal{H}_{r}(n,p)$ by the aforementioned observation); next, we show how a local algorithm on $\mathcal{H}_{r}(n,p)$ can be well-approximated by a low-degree one, completing the proof. For the first step, we adapt a local algorithm of Nie and Verstraëte on the $\Delta$-regular $r$-uniform hypertree $T_{\Delta}^{r}$ [NV21] to one on $\mathbb{T}_{d}^{GW}$ for an appropriate $\Delta\coloneqq\Delta(d)>0$. The second step follows a similar approach to that of Wein on graphs.
@@
The proof of our impossibility result (Theorem 1.6) falls into a line of work initiated by Gamarnik and Sudan [GJW20], who studied local algorithms for independent sets of $\mathcal{H}_{2}(n,p)$. The proof of their impossibility result relies on the so-called Overlap Gap Property (OGP). In particular, they show the following: if $I_{1}$ and $I_{2}$ are “large” independent sets, then either $I_{1}\cap I_{2}$ is a “large” independent set or it is “small”, i.e., either $I_{1}$ and $I_{2}$ contain a lot of common vertices or are nearly disjoint. They then show that if a local algorithm succeeds in finding a large independent set, it can be used to construct two independent sets violating the OGP condition. Rahman and Virág [RV17] improved upon their results by considering a more intricate “forbidden” structure involving many independent sets as opposed to just two. Their approach inspired further works in a number of different areas (see for example [GS14a, CHH17, Che+19]). For low-degree polynomial algorithms, Wein applied this “ensemble” variant of the OGP in the proof of his impossibility result for independent sets in $\mathcal{H}_{2}(n,p)$ [Wei22]. In our proof, we employ the ensemble-OGP in a similar fashion by considering a forbidden structure that involves many independent sets across many correlated random hypergraphs. We show that with high probability this structure does not exist, and any low-degree algorithm that can construct a large independent set can also be used to construct an instance of this structure, leading to a contradiction.
@@
Let us first discuss the statistical threshold (Theorem 1.8). The proof of the upper bound follows by a standard application of the first moment method. For the lower bound, our approach is inspired by that of Frieze in [Fri90] where he proves a high probability bound on the asymptotic size of the maximum independent set in $\mathcal{H}_{2}(n,p)$ (Krivelevich and Sudakov used a similar approach in [KS98] for $\mathcal{H}_{r}(n,p)$). Here, we fix a partition $P$ of the vertex set and we say an independent set $I$ is $P$-independent if $I$ contains at most $1$ vertex from each set in $P$. For an appropriate partition $P$, we show that there exists a $\bm{\gamma}$-balanced $P$-independent set of the desired size with high probability. A similar approach was employed by Perkins and the second author in [PW24] where they consider the case when $r=2$. For $r\geq 3$, however, certain combinatorial arguments no longer hold, requiring a more extensive analysis. The details are provided in §4.1.
The low-degree algorithm we describe for our achievability result in Theorem 1.9 is inspired by recent work of the first author who considered balanced independent sets of deterministic $r$-partite hypergraphs [Dha23]. While their primary interest was on the existence of large balanced independent sets, the proof is constructive and yields a simple degree-$1$ algorithm. The goal is to adapt this algorithm to construct $\bm{\gamma}$-balanced independent sets in $\mathcal{H}(r,n,p)$. We remark that for $r=2$, a similar approach was used by Perkins and the second author [PW24], where they describe an algorithm inspired by work of Chakraborty [Cha23]. In their setting, they note that it must be the case that $\gamma_{i}\leq 1/2$ for some $i\in\{1,2\}$. This allows them to apply Chakraborty’s procedure directly. In order for us to apply the first author’s procedure directly, we would require $\gamma_{i}\leq 1/r$ for at least $r-1$ indices $i\in[r]$ (which need not be the case). Therefore, certain modifications are necessary. We describe this modification in §4.2 describing a degree-$1$ polynomial algorithm in order to prove our achievability result.
@@ 1.5 Concluding Remarks
In this work, we consider the computational hardness of finding large (balanced) independent sets in random hypergraphs. We focus on low-degree polynomials, a powerful class of algorithms considered to be a useful proxy for computationally efficient algorithms. For the Erdős–Rényi hypergraph $\mathcal{H}_{r}(n,p)$, we determine the threshold for success of such algorithms in constructing independent sets, recovering and generalizing results of Wein [Wei22]. For the Erdős–Rényi $r$-partite hypergraph $\mathcal{H}(r,n,p)$, we prove a matching threshold for balanced independent sets, generalizing results of Perkins and the second author [PW24]. In both problems, our results indicate a statistical-computational gap of a multiplicative factor of $r^{-1/(r-1)}$ and we conjecture this to be true.
A natural extension of our results would be to consider denser hypergraphs, i.e., larger values of the probability parameter $p$. For $\mathcal{H}_{2}(n,p)$ where $p=\Theta(1)$, Matula and Karp’s results [Mat76, Kar76] indicate a statistical-computational gap of a multiplicative factor of $1/2$ for independent sets. It is not clear how to describe Karp’s algorithm as a low-degree polynomial and so the question of the low-degree threshold in this regime still remains open. Furthermore, to the best of our knowledge, hypergraph versions of these results have not yet appeared in the literature, which begs the following question: is there a statistical-computational gap of a multiplicative factor of $r^{-1/(r-1)}$ for independent sets in $\mathcal{H}_{r}(n,p)$ when $p=\Theta(1)$?
By now, OGP based methods have become standard in proving low-degree thresholds for optimization problems on graphs (see for example [Wei22, KM22, DGH23, GJW24, PW24]). Our work is the first to apply the low-degree framework and the OGP based proof strategy to optimization problems on random hypergraphs. A potential future line of inquiry would be to adapt this strategy to other optimization problems in this more general setting.
@@ 2.2 Local Algorithms on Hypergraphs
As mentioned in §1.4, we will consider two kinds of hypertrees in our proof of Theorem 1.5. First, we describe the $r$-uniform Poisson$(d)$ Galton-Watson Hypertree (denoted $\mathbb{T}_{d}^{GW}$) introduced by Pal and Zhu in [PZ21]. The process generates a rooted $r$-uniform hypertree $(T,o)$ as follows:
@@ 2.3 Interpolating Paths and Stable Algorithms
The following lemma will be crucial to proving our impossibility results. The proof follows an identical argument to that of [Wei22, Lemmas 2.7 and 2.8] on graphs and so we omit it here.
@@ 3.1.1 Local Algorithms: from $T_{\Delta}^{r}$ to $\mathbb{T}_{d}^{GW}$
We start by describing the following random greedy algorithm analyzed in [NV21], which generates an independent set in an $r$-uniform $\Delta$-regular hypergraph $H$ with large girth (which implies the case of $T_{\Delta}^{r}$):
@@
We remark that while this algorithm as described is not local, it can be simulated by one that is (this can be inferred by the discussion in [NV21, §2]). Nie and Verstraëte [NV21, Theorem 4] showed that the random greedy algorithm outputs an independent set of density $(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log\Delta}{\Delta}\right)^{1/r-1}$ on $T_{\Delta}^{r}$.
Following an idea of Rahman and Virág in [RV17], we will show that a local algorithm finding an independent set in $T_{\Delta}^{r}$ can be adapted to construct an independent set in $\mathbb{T}_{d}^{GW}$ as well. We remark that the parameter $d$ we choose will not be equal to $\Delta$, however, it will be chosen such that $\log d/d=(1\pm o_{d}(1))\log\Delta/\Delta$. In particular, the independent set output by our local algorithm on $\mathbb{T}_{d}^{GW}$ will have the same density as the one in $T_{\Delta}^{r}$ for $d$ sufficiently large. Let $E(d,\Delta)$ denote the event that the root of $\mathbb{T}_{d}^{GW}$ and all of its neighbors have degree at most $\Delta$. With this definition in hand, we prove the following proposition:
@@ 3. Step 3:
From here, we may complete the proof by an exponential moment argument to bound $p(d,\Delta-1)$. The proof follows identically to that in [RV17, Lemma 4.6] and so we omit the details. ∎
@@ 3.1.2 From Local to Low-Degree
The proof follows by combining Markov’s inequality with a hypergraph version of Proposition 12.3 [Bar+19] (for which the proof is identical, mutatis mutandis). ∎
@@
The proof follows by a hypergraph version of [Bar+19, Lemma 12.4] (for which the proof is identical, mutatis mutandis), and linearity of expectation. ∎
@@
Before we prove the main result, we make an analogous observation to the one in [Wei22, Lemma 2.11] on graphs. In particular, a random polynomial can be converted to a deterministic polynomial that works almost as well.
@@ 3.2 Forbidden Structures
In this section, we will define two “forbidden structures” in order to assist with our proof of Theorem 1.6. Before we define these structures and prove the corresponding results, we state the following theorem due to Krivelevich and Sudakov on the size of the maximum independent set in $\mathcal{H}_{r}(n,p)$ (this result wasn’t explicitly stated, but can inferred from Corollary 1 in their paper) [KS98].
@@
$\left(1-\frac{d}{{n-1\choose r-1}}\right)^{E}\leq\exp\left\{-\frac{Ed}{{n-1\choose r-1}}\right\}.$
@@
$\mathbb{E}[N]$ $=n^{2K}(T+1)^{K}\sup_{\{a_{k}\},\{b_{k}\}}\exp\left\{\frac{1}{r-1}\Phi\log d\left(a_{1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)\right\}\cdot\exp\left(-\frac{d}{{n-1\choose r-1}}E\right)$
@@
3.3 Proof of the Intractability Result in $\mathcal{H}_{r}(n,p)$

### v2

@@ 1.2 The Low-Degree Framework
We must formally define what it means for such a polynomial to find an independent set in a hypergraph. We follow the notation and definitions from [wein2022optimal], in which a rounding procedure is used to produce an independent set of $\mathcal{H}_{2}(n,p)$ based on the output $f(A)$. For brevity, we just state the following definition in this section for ordinary hypergraphs (the multipartite analog can be inferred as it is nearly identical).
@@
$V^{\eta}_{f}(A,\omega)=\begin{cases}\tilde{I}&\text{if }|I\setminus\tilde{I}|+|J|\leq\eta n;\\ \varnothing&\text{otherwise}.\end{cases}$
@@ 1.3 Main Results
$k=n\,(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{1/(r-1)},\quad\text{and}\quad\delta=\exp\left(-Cn^{1/3}\right),$
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, the following hold.
@@
    $k_{j}\leq\gamma_{j}\,r\,n(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@
    $k_{j}\,\geq\,\gamma_{j}\,r\,n(1+\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@
Setting $\gamma_{i}=1/r$ for each $i$ provides evidence toward Conjecture 2. For arbitrary $\bm{\gamma}$, we observe a statistical–computational gap of a multiplicative factor of $\gamma_{i^{\star}}^{1/(r-1)}$ for low-degree polynomial algorithms.
@@ 1.4 Proof Overview
Let us first discuss the achievability result (Theorem 1.5). It is a well-known fact that sparse random graphs are locally tree-like. In [wein2022optimal], Wein showed that the $s$-neighborhood of a vertex in $\mathcal{H}_{2}(n,d/n)$ can be well approximated by the Poisson Galton-Watson tree, a random graph model for generating trees. In [pal2021community], Pal and Zhu generalized this model to the hypergraph setting, describing the $r$-uniform Poisson Galton-Watson Hypertree $\mathbb{T}_{d}^{GW}$ (see §2.2 for a formal description). We observe that for $p=d/\binom{n-1}{r-1}$, the hypergraph $\mathcal{H}_{r}(n,p)$ is locally $\mathbb{T}_{d}^{GW}$-like, which implies that the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding algorithm evaluated at the root of $\mathbb{T}_{d}^{GW}$. Our proof now follows in two steps: first, we describe a local algorithm for finding large independent sets in $\mathbb{T}_{d}^{GW}$ (which can be adapted to one on $\mathcal{H}_{r}(n,p)$ by the aforementioned observation); next, we show how a local algorithm on $\mathcal{H}_{r}(n,p)$ can be well-approximated by a low-degree one, completing the proof. For the first step, we adapt a local algorithm of Nie and Verstraëte on the $\Delta$-regular $r$-uniform hypertree $T_{\Delta}^{r}$ [nie2021randomized] to one on $\mathbb{T}_{d}^{GW}$ for an appropriate $\Delta\coloneqq\Delta(d)>0$. The second step follows a similar approach to that of Wein on graphs.
@@
The proof of our impossibility result (Theorem 1.6) falls into a line of work initiated by Gamarnik and Sudan [gamarnik2020low], who studied local algorithms for independent sets of $\mathcal{H}_{2}(n,p)$. The proof of their impossibility result relies on the so-called Overlap Gap Property (OGP). In particular, they show the following: if $I_{1}$ and $I_{2}$ are “large” independent sets, then either $I_{1}\cap I_{2}$ is a “large” independent set or it is “small”, i.e., either $I_{1}$ and $I_{2}$ contain a lot of common vertices or are nearly disjoint. They then show that if a local algorithm succeeds in finding a large independent set, it can be used to construct two independent sets violating the OGP condition. Rahman and Virág [rahman2017local] improved upon their results by considering a more intricate “forbidden” structure involving many independent sets as opposed to just two. Their approach inspired further works in a number of different areas (see for example [gamarnik2014performance, coja2017walksat, chen2019suboptimality]). For low-degree polynomial algorithms, Wein applied this “ensemble” variant of the OGP in the proof of his impossibility result for independent sets in $\mathcal{H}_{2}(n,p)$ [wein2022optimal]. In our proof, we employ the ensemble-OGP in a similar fashion by considering a forbidden structure that involves many independent sets across many correlated random hypergraphs. We show that with high probability this structure does not exist, and any low-degree algorithm that can construct a large independent set can also be used to construct an instance of this structure, leading to a contradiction.
@@
Let us first discuss the statistical threshold (Theorem 1.8). The proof of the upper bound follows by a standard application of the first moment method. For the lower bound, our approach is inspired by that of Frieze in [frieze1990independence] where he proves a high probability bound on the asymptotic size of the maximum independent set in $\mathcal{H}_{2}(n,p)$ (Krivelevich and Sudakov used a similar approach in [krivelevich1998chromatic] for $\mathcal{H}_{r}(n,p)$). Here, we fix a partition $P$ of the vertex set and we say an independent set $I$ is $P$-independent if $I$ contains at most $1$ vertex from each set in $P$. For an appropriate partition $P$, we show that there exists a $\bm{\gamma}$-balanced $P$-independent set of the desired size with high probability. A similar approach was employed by Perkins and the second author in [perkins2024hardness] where they consider the case when $r=2$. For $r\geq 3$, however, certain combinatorial arguments no longer hold, requiring a more extensive analysis. The details are provided in §4.1.
The low-degree algorithm we describe for our achievability result in Theorem 1.9 is inspired by recent work of the first author who considered balanced independent sets of deterministic $r$-partite hypergraphs [dhawan2023balanced]. While their primary interest was on the existence of large balanced independent sets, the proof is constructive and yields a simple degree-$1$ algorithm. The goal is to adapt this algorithm to construct $\bm{\gamma}$-balanced independent sets in $\mathcal{H}(r,n,p)$. We remark that for $r=2$, a similar approach was used by Perkins and the second author [perkins2024hardness], where they describe an algorithm inspired by work of Chakraborty [chakraborti2023extremal]. In their setting, they note that it must be the case that $\gamma_{i}\leq 1/2$ for some $i\in\{1,2\}$. This allows them to apply Chakraborty’s procedure directly. In order for us to apply the first author’s procedure directly, we would require $\gamma_{i}\leq 1/r$ for at least $r-1$ indices $i\in[r]$ (which need not be the case). Therefore, certain modifications are necessary. This modification is described in detail in §4.2 and yields a degree-$1$ polynomial algorithm in order to prove our achievability result.
@@ 1.5 Concluding Remarks
In this work, we consider the computational hardness of finding large (balanced) independent sets in random hypergraphs. We focus on low-degree polynomials, a powerful class of algorithms considered to be a useful proxy for computationally efficient algorithms. For the Erdős–Rényi hypergraph $\mathcal{H}_{r}(n,p)$, we determine the threshold for success of such algorithms in constructing independent sets, recovering and generalizing results of Wein [wein2022optimal]. For the Erdős–Rényi $r$-partite hypergraph $\mathcal{H}(r,n,p)$, we prove a matching threshold for balanced independent sets, generalizing results of Perkins and the second author [perkins2024hardness]. In both problems, our results indicate a statistical–computational gap of a multiplicative factor of $r^{-1/(r-1)}$, that is, the low-degree threshold depends on $r$ for $r$-uniform hypergraphs. This is not entirely expected and cannot be easily deduced from the graph case.
A natural extension of our results would be to consider denser hypergraphs, i.e., larger values of the probability parameter $p$. For $\mathcal{H}_{2}(n,p)$ where $p=\Theta(1)$, Matula and Karp’s results [matula1976largest, karp1976probabilistic] indicate a statistical–computational gap of a multiplicative factor of $1/2$ for independent sets. It is not clear how to describe Karp’s algorithm as a low-degree polynomial and so the question of the low-degree threshold in this regime still remains open. Furthermore, to the best of our knowledge, hypergraph versions of these results have not yet appeared in the literature, which begs the following question: is there a statistical–computational gap of a multiplicative factor of $r^{-1/(r-1)}$ for independent sets in $\mathcal{H}_{r}(n,p)$ when $p=\Theta(1)$?
By now, OGP based methods have become standard in proving low-degree thresholds for optimization problems on graphs (see for example [wein2022optimal, huang2022tight, du2023algorithmic, gamarnik2024hardness, perkins2024hardness]). Our work is the first to apply the low-degree framework and the OGP based proof strategy to optimization problems on random hypergraphs. A potential future line of inquiry would be to adapt this strategy to other optimization problems in this more general setting.
@@ 2.2 Local Algorithms on Hypergraphs
As mentioned in §1.4, we will consider two kinds of hypertrees in our proof of Theorem 1.5. First, we describe the $r$-uniform Poisson$(d)$ Galton-Watson Hypertree (denoted $\mathbb{T}_{d}^{GW}$) introduced by Pal and Zhu in [pal2021community]. The process generates a rooted $r$-uniform hypertree $(T,o)$ as follows:
@@ 2.3 Interpolating Paths and Stable Algorithms
The following lemma will be crucial to proving our impossibility results. The proof follows an identical argument to that of [wein2022optimal, Lemmas 2.7 and 2.8] on graphs and so we omit it here.
@@ 3.1.1 Local Algorithms: from $T_{\Delta}^{r}$ to $\mathbb{T}_{d}^{GW}$
We start by describing the following random greedy algorithm analyzed in [nie2021randomized], which generates an independent set in an $r$-uniform $\Delta$-regular hypergraph $H$ with large girth (which implies the case of $T_{\Delta}^{r}$):
@@
We remark that while this algorithm as described is not local, it can be simulated by one that is (this can be inferred by the discussion in [nie2021randomized, §2]). Nie and Verstraëte [nie2021randomized, Theorem 4] showed that the random greedy algorithm outputs an independent set of density $(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log\Delta}{\Delta}\right)^{1/r-1}$ on $T_{\Delta}^{r}$.
Following an idea of Rahman and Virág in [rahman2017local], we will show that a local algorithm finding an independent set in $T_{\Delta}^{r}$ can be adapted to construct an independent set in $\mathbb{T}_{d}^{GW}$ as well. We remark that the parameter $d$ we choose will not be equal to $\Delta$, however, it will be chosen such that $\log d/d=(1\pm o_{d}(1))\log\Delta/\Delta$. In particular, the independent set output by our local algorithm on $\mathbb{T}_{d}^{GW}$ will have the same density as the one in $T_{\Delta}^{r}$ for $d$ sufficiently large. Let $E(d,\Delta)$ denote the event that the root of $\mathbb{T}_{d}^{GW}$ and all of its neighbors have degree at most $\Delta$. With this definition in hand, we prove the following proposition:
@@ 3. Step 3:
From here, we may complete the proof by an exponential moment argument to bound $p(d,\Delta-1)$. The proof follows identically to that in [rahman2017local, Lemma 4.6] and so we omit the details. ∎
@@ 3.1.2 From Local to Low-Degree
The proof follows by combining Markov’s inequality with a hypergraph version of Proposition 12.3 [barak2019nearly] (for which the proof is identical, mutatis mutandis). ∎
@@
The proof follows by a hypergraph version of [barak2019nearly, Lemma 12.4] (for which the proof is identical, mutatis mutandis), and linearity of expectation. ∎
@@
Before we prove the main result, we make an analogous observation to the one in [wein2022optimal, Lemma 2.11] on graphs. In particular, a random polynomial can be converted to a deterministic polynomial that works almost as well.
@@ 3.2 Forbidden Structures
In this section, we will define two “forbidden structures” in order to assist with our proof of Theorem 1.6. Before we define these structures and prove the corresponding results, we state the following theorem due to Krivelevich and Sudakov on the size of the maximum independent set in $\mathcal{H}_{r}(n,p)$ (this result wasn’t explicitly stated, but can inferred from Corollary 1 in their paper) [krivelevich1998chromatic].
@@
$\left(1-\frac{d}{{n-1\choose r-1}}\right)^{E}\leq\exp\left(-\frac{Ed}{{n-1\choose r-1}}\right).$
@@
$\mathbb{E}[N]$ $=n^{2K}(T+1)^{K}\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\frac{1}{r-1}\Phi\log d\left(a_{1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)\right)\cdot\exp\left(-\frac{d}{{n-1\choose r-1}}E\right)$
@@
3.3 Proof of the Intractability Result